OpenStudy (anonymous):

inverse function of f(x)=sqrt(x^2+5x) An explanation of method would be great!

OpenStudy (anonymous):

y = sqrt(x^2 + 5x) y^2 = x^2 + 5x y^2 = x^2 + 5x + 2.5^2 - 2.5^2 y^2 = (x+2.5)^2 - 2.5^2 sqrt(y^2 + 2.5^2) = x +2.5 x = sqrt(y^2 + 2.5^2) - 2.5 so y = sqrt(x^2 + 2.5^2) - 2.5 is the inverse :)

OpenStudy (anonymous):

Thanks!